### v1

@@ 5.2 Non-affine Symmetry Enforcement
Our second experiment compares ordinary polynomial regression with polynomial regression utilizing our symmetry enforcement method. First, we generate 2000 datapoints $(x_{i},y_{i})$ in the plane in a “thin strip” configuration: each $x_{i}$ is generated from $\text{Uniform}(-4,4)$, and $y_{i}=0+\epsilon_{i}$, where $\epsilon_{i}\in(-0.1,0.1)$ is a random noise component. For each point, we assign regression targets according to the function
@@
To apply “ordinary” polynomial regression, we seek to minimize the MSE between the output of the function $f$ given in equation 26 and the output of an arbitrary degree four polynomial, iteratively updating the coefficients. We use gradient descent, a learning rate of $10^{-4}$, and $100,000$ epochs. [^5] We ignore the least squares exact solution, since this method is generally unavailable in machine learning. This results in a training MSE of $0.0159$ and a learned function given approximately as
@@
$\cos(\theta_{\text{train}})=\dfrac{\langle\hat{f},f\rangle_{\text{train}}}{||\hat{f}||_{\text{train}}||f||_{\text{train}}}\approx 0.9999.$
[…]
$\cos(\theta_{\text{test}})=\dfrac{\langle\hat{f},f\rangle_{\text{test}}}{||\hat{f}||_{\text{test}}||f||_{\text{test}}}\approx 0.8847.$
Additionally, a visualization of the predictions for $\hat{f}$ are given in Figure 1. To estimate $f$ using symmetry enforcement, we apply equation 17 with $\lambda(t)=0.5$, with both loss functions being the MSE. Training, as before, for $100,000$ epochs, we obtain a (total) training loss function minimum of $1.4\cdot 10^{-6}$, with
(29) $\hat{f}_{s}\approx 0.0001+0.0004y^{2}+2.0000x^{4}-2.0000x^{2}y^{2}+0.9594y^{4}.$
We find that
$\cos(\theta_{\text{train}})\approx 1.0000,\qquad\cos(\theta_{\text{test}})\approx 0.9998,$
and, as Figure 1 reveals, model predictions whose symmetries appear to be more in line with that of the ground truth function.
@@ 5.3 Isometry Discovery
We consider a 3-dimensional classification problem in which points $(x,y,z)$ are assigned to class 1 if $f(x,y,z)>0$ and to class 0 otherwise, where
$f(x,y,z)=x^{2}+\tfrac{y^{2}}{2}-yz+\tfrac{z^{2}}{2}-\tfrac{1}{2}.$
In order to learn the class separation, we map the original data to a 6-dimensional Euclidean feature space only using the degree-2 terms of a polynomial feature map:
@@
Next, a logistic regression classifier is trained in the 6-d feature space, producing a decision boundary that, when translated back to the original coordinates, approximately corresponds to the quadratic form $f(x,y,z)=0$. We wish to learn the isometries of the model probabilities: that is, the Killing vectors which annihilate
[…]
The metric $g$ admits three linearly independent Killing vectors, the flows of which are 3-d rotations:
@@
We now construct an arbitrary linear combination
$\hat{X}=aX_{1}+bX_{2}+cX_{3}$
and optimize $(a,b,c)$ via equation (14). We note that a ground truth solution $X$ has coefficients $(0,-1,1)$. Using the L1 Loss function and the (Riemannian) Adagrad optimizer with a learning rate of $0.01$, we obtain the coefficients $(0.0055,0.7076,-0.7066)$, and the similarity between $X$ and $\hat{X}$ is estimated to be $0.99992$.
@@ 6 Conclusion
While most current state of the art experiments focus, in the context of continuous symmetry detection, on affine symmetries, we have outlined a method to estimate symmetries including and beyond affine transformations using vector fields. Our method relies on estimating machine learning functions such as a probability distribution, continuous targets for regression, or functions which characterize an embedded submanifold (level set estimation). Vector fields are constructed as arbitrary linear combinations of a chosen basis of vector fields, and the parameters that are the coefficients of the linear combination are optimized, often using constrained optimization. Herein, we have typically expressed the coefficients of our vector fields in terms of polynomial functions.
When searching for affine symmetries in low dimensions, our method is appealing when compared with many current methods. This is due primarily to the relative ease of implementation. Additionally, when compared with the state of the art, our method appears to offer a computational advantage while at least competing and often outperforming in terms of accuracy, especially for more complex symmetries than affine ones [23].
Our infinitesimal generators are a step removed from the symmetry group than what is typical of symmetry detection methods. We have proposed building invariant architectures based on the identification of features which are invariant with respect to the discovered vector fields, resulting in an invariant model, since the information fed to the model is approximately invariant with respect to the symmetry group. We also proposed a means of symmetry enforcement in the training of a model though symmetry regularization.

### v2

@@ 6.2 Non-affine Symmetry Enforcement
Our second experiment compares ordinary polynomial regression with polynomial regression using our symmetry enforcement method. First, we generate 2000 datapoints $(x_{i},y_{i})$ in the plane in a “thin strip” configuration: each $x_{i}$ is generated from $\text{Uniform}(-4,4)$, and $y_{i}=0+\epsilon_{i}$, where $\epsilon_{i}\in(-0.1,0.1)$ (distributed uniformly) is a random noise component. For each point, we assign regression targets according to the function
@@
To apply “ordinary” polynomial regression, we seek to minimize the MSE between the output of the function $f$ given in Equation (20) and the output of an arbitrary degree four polynomial, iteratively updating the coefficients. We use gradient descent, a learning rate of $10^{-4}$, and $100,000$ epochs.This results in a training MSE of $0.0159$ and a learned function given approximately as
@@ 6.3 Isometry Discovery
We consider a 3-dimensional classification problem in which points $(x,y,z)$ are assigned to class 1 if $f(x,y,z)>0$ and to class 0 otherwise, where $f(x,y,z)=x^{2}+\tfrac{y^{2}}{2}-yz+\tfrac{z^{2}}{2}-\tfrac{1}{2}.$ To learn the class separation, we map the original data to a 6-dimensional Euclidean feature space only using the degree-2 terms of a polynomial feature map:
@@
Next, a logistic regression classifier is trained in the 6-d feature space, producing a decision boundary that, when translated back to the original coordinates, approximately corresponds to the quadratic form $f(x,y,z)=0$. We wish to learn the isometries of the model probabilities, that is, the Killing vectors which annihilate $p(x,y,z)\;=\;\frac{1}{\,1+e^{-\,f(x,y,z)}\,}.$ The metric $g$ admits three linearly independent Killing vectors, the flows of which are 3-d rotations:
@@
We now construct an arbitrary linear combination $\hat{X}=aX_{1}+bX_{2}+cX_{3}$ and optimize $(a,b,c)$ via Equation ((12)). We note that a ground truth solution $X$ has coefficients (0,1,1)(0,-1,1). Using the L1 Loss function and the (Riemannian) Adagrad optimizer with a learning rate of $0.01$, we obtain the coefficients (0.0055,0.7076,0.7066)(0.0055,0.7076,-0.7066), and the similarity between $X$ and $\hat{X}$ is estimated to be $0.99992$.
@@ 7 Conclusion
While most current state of the art experiments focus, in the context of continuous symmetry detection, on affine symmetries, we have outlined a method to estimate symmetries including and beyond affine transformations using vector fields. Our method takes as input an estimated machine learning function such as a probability distribution, continuous targets for regression, or functions which characterize an embedded submanifold (obtained via level set estimation). Vector fields are constructed as arbitrary linear combinations of a chosen basis of vector fields, and the parameters that are the coefficients of the linear combination are optimized, often using constrained optimization. Herein, we have typically expressed the coefficients of our vector fields in terms of polynomial functions, excepting isometries, in which a linear combination of Killing vectors is formed. Future work includes detecting other types of symmetries, including angle-preserving transformations whose infinitesimal generators are known as conformal Killing vectors [Shaw2].
When searching for affine symmetries in low dimensions, our method is appealing when compared with many current methods. This is due primarily to the relative ease of implementation. Additionally, when compared with the state of the art, our method offers a computational advantage while at least competing and often outperforming in terms of accuracy, especially for more complex symmetries than affine ones [Shaw1].
Infinitesimal generators are a step removed from the group action. We have proposed building invariant architectures based on the identification of features which are invariant with respect to the discovered vector fields, resulting in an invariant model, since the information fed to the model is approximately invariant with respect to the group action. We also proposed a means of symmetry enforcement in the training of a model though symmetry regularization. Both methods may help to improve model generalization.